Previous

Calendar

July

2008

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

29

JD181

30

JD182

1

JD183

2

JD184
SpB Returns

3

JD185
Site Visit (JB, SpB, SH, NT, JH, SL): Details , Photos (34)

span @ 610 psi, N2 @ 675 psi

from 10:15-10:45 MST, changed domes on 26m REBS Rnet sensor

downloaded from cnr, prof, nctc, met, soil, tcs, ucb and uc loggers

changed internal chemicals in the LI-7500 (sn 75H-0084) on the main tower

swapped 15 flasks at the robot

at ~10 MST, increased flow rate in prof system from "9" to "14"

at ~13:00 MST, prof pump (UN828 KNI) at the base of the main tower was dead, we replaced it with a new pump

Cloudy down on the plains; sunny at the tower



4

JD186
CU Holiday (Independence Day)

5

JD187

6

JD188
Site Visit (SO, LL): Details

2.5m csat was moved (by accident). Returned to approx. original position.

7

JD189

8

JD190
Site Visit (JB, SH, NT, CB, SL): Details

prof "flow-meter" ball was at ~13, increased the flow to 15

between 13-14 MST, chan 204 (met logger) was dropping data points (seems ok now)

HEAVY hail/rain storm 12:15-12:45 MDT

9

JD191

10

JD192

11

JD193
Site Visit (SpB, SO, LL, M, RM, SH, SL): Details , Photos (11)

at ~11:00 MST, added KZ PAR-LITE sensor as PAR2.up to CNR logger

started install of MATLAB on computers for the flux class.


12

JD194

13

JD195
Site Visit (SpB, RM, DM, etc): Details

Met with campbell folks; finished MATLAB install at MRS.


14

JD196
Site Visit (JB, SpB, SO, LL, Mike): Details , Photos (18)

swapped 15 flasks at the west robot

brought the N2 cylinder (CC193097) from Como Creek out to the flux tower, and brought a new span cylinder (CC282376) from the trailer to the tower

downloaded from the soil and tcs loggers

from 10:30-11:15 MST, replaced the ec gast pump

at 11:30 MST, cleaned the krypton optics

at 11:45 MST, connected power to fan for 2.5m REBs Rnet sensor

helped SO at tram in the afternoon

Start of "Flux" class at MRS.

15

JD197
Site Visit (SpB, MP/SH/CB, NT/SL, JM): Details

16

JD198
Site Visit (SpB, RM, MP/SH/CB, NT/SL, etc): Details , Photos (7)

from 9-noon MDT, Campbell Sci instrumentation demonstration for Flux class (at trailer).

from 11:00-11:30 MST, Replaced N2 at tower (New cylinder= CC193097 at 1900 psi, Old cylinder= CC262308, at 150 psi)

17

JD199
Site Visit (SpB, RM, DB, JH, MP/SH/CB, NT/SL, etc): Details , Photos (6)

from 9-noon MDT, Tower visit by the Flux class.

JB Off

Cool and Overcast--a brief rain shower in the afternoon

18

JD200
Site Visit (SpB, RM, DB, SO, LL): Details , Photos (7)

from 11:32-11:48 MST, Changed profiler (li-6251) dessicant

lowered flow rates of N2/Span gas for LI-6262 (sample flow rate is a bit low)

tram work (afternoon)

better ventilation system added to bowling shed (DB)

19

JD201

20

JD202

21

JD203
Site Visit (SpB, MP/SH/CB/RaM, NT/SL): Details , Photos (11)

from 11:38-12:08 MST, replace 2-micron Nupro filters in EC/LI-6262 system.

re-adjusted flow rates of N2/Span gas for LI-6262 (since sample flow rate is now higher)

swapped 15 flasks at the west robot

Clear skies (and HOT) in morning--clouded up and cooled off in the afternoon

22

JD204
Short (1-second?) Power outage at 12:53:51 MDT

23

JD205
Site Visit (SpB, MP/JH/CB, NT/SL): Details , Photos (5)

worked on PAR setup (swapped card in grove 1, and then moved ladder to grove 2)

Overcast in morning--a brief (fairly hard) rain shower in the afternoon

24

JD206

25

JD207
End of the "Flux" class at MRS.

JB returns to work

26

JD208

27

JD209

28

JD210
Site Visit (SpB/JB, MP/SH/BH/NM, NT/SL): Details , Photos (11)

worked on PAR setup (worked on two trees in grove 2)

swapped 15 flasks at the west robot

Cool and overcast in morning--some rain and a bit of lightning/thunder in afternoon

29

JD211

30

JD212
High Ozone alerts issued for Denver-metro and Front Range region

31

JD213

1

JD214
Clear skies (and HOT) all day today

JB with Merced

2

JD215